LBC Technical Observing/Restart after shutdown:
UT 2011 Sep 25

Observer: D.Thompson
Telescope Operator: D.Gonzalez Huerta
Instrument Support: J.Morris
Telescope Support:
SW support:

UT 11:40 Still closed for thick clouds.

-- DavidThompson - 25 Sep 2011
Topic revision: r1 - 25 Sep 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback